NERSCPowering Scientific Discovery Since 1974

May 10

Attending:  Eric, Jay, Katie, Andrei, Jeff P., Marjorie

Utilization:  Cluster has been completely full.

- Alicernc (alicesgm?) has been transferring on batch nodes.

Outages:  none

Upcoming Downtimes:  nothing until post-QM.  Still plan to switch moe and upgrade GPFS on elizas.

Procurements/New Hardware:  We will be getting ~1 rack's worth of (ALICE) comput nodes in the next month, ~500 cores.  2007 nodes are out of warranty - this is racks 9+10, ~40 noes or 300 cores.  These will go away.

Training:  Discussed the possiblity of Eric hosting a training session for PDSF users.  A preference was expressed for web documentation instead.

New batch system:  All is working except io resources.  However, the additional benefits and feature might outweigh this deficiency and we will press on with the transition.

Grid Issues:  Andrei reports success with ATLAS grid transfers - however, files were much smaller than when there were problems before.

- Discussed disk usage monitoring and Jay will look into generating data for the different groups.

- Discuss exit status 1 jobs:  This is not understood at present...